One has the following information regarding a synthetic account: Depreciation: 2.4 times larger than the subsidies Indirect Tax on goods and imports: 17.5 billion Euro Gross domestic product (market prices): 12.1 times larger than the indirect tax on goods and imports Subsidies: 4.3% from the Gross domestic product (market prices) Using a synthetic account please compute the value of the Net domestic product!

with steps of solutions

a.The value of Net Domestic Product is 181.5 billion Euro

b.The value of Net Domestic Product is 133.7 billion Euro

c.The value of Net Domestic Product is 143.0 billion Euro

d.The value of Net Domestic Product is 164.2 billion Euro

e.The value of Net Domestic Product is 169.5 billion Euro


1
Expert's answer
2022-01-27T07:53:31-0500

NNP = GNP - A


"GDP=17.5\times12.1=211.75"


"Subsidies=0.043\times211.75=9.10525"


"A=2.4\times9.10525=21.8526"


"GNP=GDP+17.5-9.10525=211.75+17.5-9.10525=220.14475"


"NNP=220.14475-21.8526=198.29215"